OpenStudy (anonymous):

Which of the following is the solution to the equation 8^(y + 2) = 2^y ?

OpenStudy (rsadhvika):

write 8 as 2^3

OpenStudy (rsadhvika):

2^3(y+2) = 2^y

OpenStudy (rsadhvika):

since bases are equal, exponents also will be equal

OpenStudy (rsadhvika):

3(y+2) = y

OpenStudy (rsadhvika):

solve y

OpenStudy (anonymous):

help me.. theres two ys?

OpenStudy (anonymous):

@rsadhvika

OpenStudy (rsadhvika):

3(y+2) = y 3y + 6 = y 2y = -6 y = -3

OpenStudy (anonymous):

oh okay. thank you.

OpenStudy (rsadhvika):

yw